I am not even sure where to start with this problem. Any assistance would be helpful. Thank you!
A survey of 12 people at a bookstore showed that five had purchased a hardcover book. If seven people are selected at random, what is the probability that four of them had purchased a hardcover book?

Answer by stanbon(48568)

A survey of 12 people at a bookstore showed that five had purchased a hardcover book. If seven people are selected at random, what is the probability that four of them had purchased a hardcover book?
-----------------------------------------------
From the survey you have P(person buys a hardcover book) = 5/12
----------------------------------
In a random sample of 7 you have a binomial problem with p = 5/12
---------------------
P(4 buy out of 7) = 7C4(5/12)^4(7/12)^3 = 0.2094
-----------------------
If you have a TI calculator that is binompdf(7,(5/12),4)
==============================
Cheers,
Stan H.